Use your calculator to find the values of the following trigonometric function to four decimal places. [8 pts.] Hint: Your calculator should be in radian mode. (radian)

a. sin38°

b. cos21°

c. tan32.7°

d. sin 200°

e. sin 6 p

f. tan 4 p 5

g. cos ( p /18)

h. sin 3 p/10
Answer #1

@ Sin (380) 0.296 3 6 -0.9963.. b) C06 (210) -0.54 7729 -0.5477 e(c) tan(39.7) = 3,391586 foong 53:3915 d) Sin 200°)=-0.83997. =-0.8.7329 le) sin (6p) = sin (6 TT) 0. (f) tan up 5 = tan(ut/